OpenStudy (loser66):

Need help, please Question 2: The women of a locality plan to have a Lunch Party as a part of celebration of Women's Day. The women plan to share the cooking work, by each preparing a dish. There were 5 options given for the Main Course. What is the probability that three friends. Ann, Mary and Jane offering to cook three different main course items. Solution: P(Ann choosing a different item) = 5/5 =1 P(Mary choosing item different from what Ann has chosen) = 4/5 P(Jane choosing an item different from both Ann's and Mary's choices ) = 3/5

OpenStudy (loser66):

I don't get why P(A) =1

satellite73 (satellite73):

because she is the first to choose, so it has to be different from the others

OpenStudy (mantar0078):

P(A)=1 because the fraction is 5/5 which would simplify to 1

OpenStudy (loser66):

why not 5C3?

OpenStudy (loser66):

why is 5/5?

satellite73 (satellite73):

forget the 5/5 that is silly

OpenStudy (mantar0078):

Same with 9/9=1 or why when you have a fraction like 6/5 you would simplify it to 1 1/5 because it is larger than 5

OpenStudy (loser66):

@mantar0078 I know it. @satellite73 cannot "forget it", because, later on, P (Mary...) =4/5 and P(Jane...) =3/5 those fractions MUST has some meaning

OpenStudy (loser66):

*have

satellite73 (satellite73):

ok sure they mean something

OpenStudy (mantar0078):

The fractions are the probability

satellite73 (satellite73):

the first person to choose has to choose something that has not been chosen before

satellite73 (satellite73):

so the probability that they choose something un-chosen is one

OpenStudy (loser66):

got that part

satellite73 (satellite73):

oh,i thought that is what you were asking

satellite73 (satellite73):

i quote "I don't get why P(A) =1"

OpenStudy (loser66):

how about 4/5, and 3/5?

satellite73 (satellite73):

ok so one item is chosen yes ?

OpenStudy (loser66):

Yes

satellite73 (satellite73):

that leaves 4 new ones and one old one pick one of the five at random, the probability you get one of the four unchosen ones is four out of five

OpenStudy (loser66):

Got you. Thank you so much

satellite73 (satellite73):

you're quite welcome